Let 𝑓 be a function which is everywhere differentiable and for which 𝑓(2) = βˆ’3 and 𝑓 β€² (π‘₯) = √π‘₯ 2 + 5. Given that 𝑔 is defined such that 𝑔(π‘₯) = π‘₯ 2𝑓 ( π‘₯ π‘₯ βˆ’ 1 ), show that 𝑔 β€² (2) = βˆ’24.


1
Expert's answer
2021-09-28T00:00:52-0400

Given that "f'(x)=\sqrt{x^2+5}"


"g(x)=x^2(f(x)-1)"

Differentiating "g(x)" w.r.t ."x," we get:

"g'(x)=2x(f(x)-1)+x^2(f'(x))"

Put "x=2"

"g'(2)=4(f(2)-1)+2^2(f'(2))"

"=4((-3)-1)+4f'(2)\\
=-16+4(3)\\
=-4"